# A Data Efficient Framework for Learning Local Heuristics

### v1

@@ Abstract
With the advent of machine learning, there have been several recent attempts to learn effective and generalizable heuristics. Local Heuristic A* (LoHA*) is one recent method that instead of learning the entire heuristic estimate, learns a “local” residual heuristic that estimates the cost to escape a region (Veerapaneni et al 2023). LoHA*, like other supervised learning methods, collects a dataset of target values by querying an oracle on many planning problems (in this case, local planning problems). This data collection process can become slow as the size of the local region increases or if the domain requires expensive collision checks. Our main insight is that when an A* search solves a start-goal planning problem it inherently ends up solving multiple local planning problems. We exploit this observation to propose an efficient data collection framework that does $<$1/10th the amount of work (measured by expansions) to collect the same amount of data in comparison to baselines. This idea also enables us to run LoHA* in an online manner where we can iteratively collect data and improve our model while solving relevant start-goal tasks. We demonstrate the performance of our data collection and online framework on a 4D $(x,y,\theta,v)$ navigation domain.
@@ 1 Introduction
Recently, several works have proposed using machine learning to obtain better heuristics (or priorities) to speed up search. We focus on methods using supervised learning which requires collecting a training dataset of optimal solutions generated by an oracle search method. For example, Kim and An (2020) uses A* to generate a training database of optimal solution values while Takahashi et al. (2021) uses a backward Dijkstra. Kaur, Chatterjee, and Likhachev (2021) learns an “expansion delay” heuristic (a proxy measure for the size of local minima) and gathers training data by using an oracle A* on planning problems and recording the expansion number for each state on the optimal path. SAIL (Bhardwaj, Choudhury, and Scherer 2017) learns a priority function by utilizing optimal cost-to-go values collected by running a backward Dijkstra oracle deployed solely for data collection.
@@
An important byproduct of this data collection mechanism is that we no longer need a separate data collection phase and can collect data when running LoHA*. Since LoHA* runs a search method during test time, we can collect local heuristic data online while using LoHA* itself. This enables us to rapidly learn from our experiences in a way not possible with other data collection techniques.
@@ 2 Preliminaries
Computing $h_{k}(s)$ requires identifying the best border state $s^{\prime}\in LRB(s)$ that minimizes $h_{gk}(s)$. This is achieved by running a local (multi-goal) A* search from $s$ using all the border states as goals with $h_{g}$ as the heuristic ($h_{g}$ estimates the cost to $s_{g}$ and not the border states). In the most common case, where the goal is not located within $LR(s)$, the search will terminate upon expanding the first state in $LRB(s)$. This heuristic residual is effective but is too slow to be computed during test time, so LoHA* approximates it with a neural network.
@@
Figure 2: A simplified example of a global A* while collecting data for local regions with $K=3$, i.e. for $s=(x,y)$ we want to reach a state $(x^{\prime},y^{\prime})$ with $x^{\prime}$ or $y^{\prime}$ 3 away. Each circle (i) depicts the $i^{th}$ state expanded by A* with arrows denoting successor $\rightarrow$ parent. Expanding (2) and (3) leads to incomplete data $D_{ic}$ for $s_{1},s_{2}$ as we have made some progress but have not reached a $LRB$ yet. Expanding $s_{4}$ and backtracking reveals that $s_{4}\in LRB(s_{1})$; we have found $LH(s_{1})$ and add it to our complete dataset $D_{c}$. We continue to expand a node (purple), backtrack (blue arrows) to update values of ancestors (blue) whose $LH(s^{\prime})$ have not been computed. Bottom right: After 10 expansions, we collect 4 complete (green) and 2 partial (blue) LH values, while we cannot collect data for leaf nodes (red).
@@ 3.2 Collecting Partial Data
Although this logic works, the amount of data collected per iteration of global search was lower than expected. This can be attributed to the fact that a large number of nodes $s$ in the global search never had an $s^{\prime}\in LRB(s)$ expanded. Instead, many nodes had made partial progress, i.e., nodes in the local region but not in the border were expanded. Visually in Figure 2 this corresponds to $s_{6}$ and $s_{7}$ which both have partial progress (via $s_{10}$) but don’t have a node in the $LRB$ expanded. Instead of completely ignoring these data points, we developed a mechanism to extract useful approximations on $h_{k}(s)$ and include it in the dataset.
@@
[…]
Figure 3: (a) compares collecting data via a ground truth oracle (Local A*) against our “Complete” and “Incomplete” collection methods. (b) plots the “Speed-up” (as measured in nodes expanded) of LoHA* trained on datasets gathered from Local A* calls (true oracle) or our incomplete data gathered from running a global A* on start-goal problems. (c) We run DE-LoHA* by solving start-goal problems, collecting data, and retraining every 5 problems. We see that DE-LoHA* can improve performance from just solving start-goal problems without needing an external oracle after the initial 5 problems.
@@
Closed List as an Obstacle Astute readers may have noticed a possible issue with directly using Theorem 1; although the states sorted identically, we have ignored the effect of the closed list of previously expanded states in the global A*. These states $\in LR(s)$ would not be re-expanded when explored through $s$ in the global search and we can thus view states in the closed list as obstacles. Note this observation has been exploited in previous heuristic search work on grid worlds (Felner, Shperberg, and Buzhish 2021). Interestingly, we found that incorporating the closed list led to no meaningful performance difference so we ignore it.
@@ 4 Experimental Results
Data Efficiency Table 3 shows how many nodes expanded are required to collect a single training data point as a function of $K$. We see that running local searches scales poorly and requires 100s of expansions for a single data point for $K>4$. We observe that as $K$ increases, the work for gathering “Complete” via backtracking increases but then saturates as when $K$ is sufficiently high, the global A* mainly adds data points on states on the solution path and few else (as the search rarely fully explores a local minima not on the path). On the flip side, incomplete remains constant regardless of $K$ as all states with any successor expanded becomes a data point (as we have partial data about its $LH(s)$ value), including states that are in local minima.
@@
One concern with using incomplete data by backtracking is that it is approximate and noisy in comparison to the accurate residuals computed by local A*. In order to understand the impact of this quality difference, in Fig 3 we plot the performance of LoHA* when using the incomplete data and as a function of the amount of nodes expanded to collect each dataset. The y-axis “Speed-up” is the multiplicative reduction of nodes expanded to find paths across 100 problems between weighted LoHA* and weighted A* with $w=4$ (e.g. 3 represents LoHA* expanding 1/3 the nodes). We see that even though incomplete data contains approximations, we can still effectively use it to learn a heuristic that leads to performance gain in substantially less data. Using a local A* requires a magnitude more work to get similar performance. We ran an ablation removing downweighing (i.e. not weighting by $\alpha(s,s^{\prime})$ and found that this reduced performance from 3.9x to 2.4x). Note that although LoHA* expands fewer nodes, the model inference introduces a large overhead which results in it taking roughly 2.3 seconds per problem compared to the baseline which takes 0.1 seconds.
@@
Online Performance Lastly, we evaluate how using our data-efficient framework DE-LoHA* enables learning local heuristics online. We initially start with a small dataset collected by solving 5 start-goal planning problems using global A* (and the backtracking logic). This small dataset is then used to train an initial local heuristic model. The learnt model is then used with DE-LoHA* to solve another 5 start-goal problems. The data accumulated while solving these problems is accumulated and then used to retrain and improve the model further (the model is retrained after solving every 5 problems). We evaluate online DE-LoHA* run with different $w$ with the corresponding weighted A* baseline. We plot the performance on 50 problems (plotting on just the 5 problems encountered is too noisy). We see that within 20 encountered start-goal problems (so 3 iterations of running DE-LoHA*), we start to get a non-trivial performance increase. As we continue solving problems and gathering data, improvement either increases ($w=16$) or saturates ($w=4,8$). This highlights how DE-LoHA* can perform well from just solving start-goal problems.
@@ 5 Conclusion
Extensions Our data efficient technique can be applied with no or small modifications to other learning methods which requires an oracle search method. When learning a cost-to-go heuristic $h(s,s^{\prime})$, existing work just uses states on the solution path with the oracle’s optimal $c(s,s_{goal})$ (Jabbari Arfaee, Zilles, and Holte 2011; Kim and An 2020). We observe how any state $s_{i}$ and an ancestor $s_{i}^{\prime}$ in the oracle’s search tree results in a valid optimal $c(s_{i}^{\prime},s_{i})$ data point that can be used as well. We can thus backtrack from $s_{i}$ to ancestors to collect data. Similarly for learning an expansion delay (Kaur, Chatterjee, and Likhachev 2021), the expansion delays between states $s_{i}^{\prime},s_{i}$ can be used. We hope that future work builds on this data efficient framework to decrease computational burden and enable online learning. Acknowledgements This material is partially supported by NSF Grant IIS-2328671.
@@ 1 Procedure Plan($S,E,s_{start},s_{goal},K$)
3             $s_{min}$ = OPEN.min() if $s_{min}==s_{goal}$ then
@@
5            Successors = Expand($s_{min}$) $\forall s^{\prime}\in$ Successors, $s^{\prime}$.parent = $s_{min}$triangleright Required anyway to backtrack solution path BacktrackDataCollection($s_{min}$) $\forall s^{\prime}\in$ Successors, OPEN.insert($s^{\prime}$)
@@ 7 Procedure BacktrackDataCollection($s$)
8       $s_{cur}$ = $s$.parent while $s_{cur}$ != null and $s_{cur}\notin$ CompletedData do
@@
10                   CompletedData[$s_{cur}$] = b ($s_{cur},s$) Remove $s_{cur}$ from IncompleteData
@@ 12                   IncompleteData[$s_{cur}$] = b ($s_{cur},s$)
14 Procedure b($s$, $s^{\prime}$)
15       $c(s,s^{\prime})=s^{\prime}.cost-s.cost$ $h_{k}=c(s,s^{\prime})+h_{g}(s^{\prime})-h_{g}(s)$ return $h_{k}$

### v2

@@ Abstract
With the advent of machine learning, there have been several recent attempts to learn effective and generalizable heuristics. Local Heuristic A* (LoHA*) is one recent method that instead of learning the entire heuristic estimate, learns a “local” residual heuristic that estimates the cost to escape a region. LoHA*, like other supervised learning methods, collects a dataset of target values by querying an oracle on many planning problems (in this case, local planning problems). This data collection process can become slow as the size of the local region increases or if the domain requires expensive collision checks. Our main insight is that when an A* search solves a start-goal planning problem it inherently ends up solving multiple local planning problems. We exploit this observation to propose an efficient data collection framework that does $<$1/10th the amount of work (measured by expansions) to collect the same amount of data in comparison to baselines. This idea also enables us to run LoHA* in an online manner where we can iteratively collect data and improve our model while solving relevant start-goal tasks. We demonstrate the performance of our data collection and online framework on a 4D $(x,y,\theta,v)$ navigation domain.
@@ 1 Introduction
Recently, several works have proposed using machine learning to obtain better heuristics (or priorities) to speed up search. We focus on methods using supervised learning which requires collecting a training dataset of optimal solutions generated by an oracle search method. For example, Kim and An (2020) uses A* to generate a training dataset of optimal solution values while Takahashi et al. (2021) uses a backward Dijkstra. Kaur, Chatterjee, and Likhachev (2021) learns an “expansion delay” heuristic (a proxy measure for the size of local minima) and gathers training data by using an oracle A* on problems and recording the expansion number for each state on the optimal path. SAIL (Bhardwaj, Choudhury, and Scherer 2017) learns a priority function by utilizing optimal cost-to-go values collected via a backward Dijkstra oracle deployed solely for data collection.
@@
An important byproduct of this data collection mechanism is that we no longer need a separate data collection phase and can collect data when running LoHA*. Since LoHA* runs a search method during test time, we can collect local heuristic data online while using LoHA* itself. This enables us to rapidly learn from our experiences in a way not possible with other data collection techniques. See https://arxiv.org/abs/2404.06728 for appendix.
@@ 2 Preliminaries
Computing $h_{k}(s)$ requires identifying the best border state $s^{\prime}\in LRB(s)$ that minimizes $h_{gk}(s)$. This is achieved by running a local (multi-goal) A* search from $s$ using all the border states as goals with $h_{g}$ as the heuristic ($h_{g}$ estimates the cost to $s_{g}$ and not the border states). In the most common case, where the goal is not located within $LR(s)$, the search terminates upon expanding the first state in $LRB(s)$. This heuristic residual is effective but slow to compute, so LoHA* approximates it with a neural network.
@@
Figure 2: A simplified example of global A* collecting data for local regions with $K=3$, i.e. for $s=(x,y)$ we want to reach a state $(x^{\prime},y^{\prime})$ with $x^{\prime}$ or $y^{\prime}$ 3 away. Each (i) depicts the $i^{th}$ state expanded with successor $\rightarrow$ parent denoted. Expanding (2), (3) leads to incomplete data $D_{ic}$ for $s_{1},s_{2}$ as we made some progress but did not reach a $LRB$ yet. Expanding $s_{4}$ and backtracking reveals that $s_{4}\in LRB(s_{1})$; we have found $LH(s_{1})$ and add it to our complete dataset $D_{c}$. We continue to expand a node (purple), backtrack (blue arrows) to update values of ancestors (blue) whose $LH(s^{\prime})$ have not been computed. Bottom right: After 10 expansions, we collect 4 complete (green) and 2 partial (blue) LH values, and cannot collect data for leaf nodes (red).
@@ 3.2 Collecting Partial Data
Although this logic works, the amount of data collected per iteration of global search was lower than expected. This was due to many nodes $s$ in the global search never having an $s^{\prime}\in LRB(s)$ expanded. Instead, many nodes had made partial progress, i.e. nodes in the local region but not in the border were expanded. In Figure 2 this corresponds to $s_{6}$ and $s_{7}$ which both have partial progress (via $s_{10}$) but don’t have a node in the $LRB$ expanded. Instead of completely ignoring these data points, we developed a mechanism to extract useful approximations on $h_{k}(s)$ and include it in the dataset.
@@
(a) Amount of work (# expansions) required to collect a single data point as $K$ increases. Local A* uses an oracle A* to produce individual data points while Complete and Incomplete gather data across global start-goal problems using our backtracking methodology.
(b) Performance vs Work for Dataset
(c) Online performance of DE-LoHA*
Figure 3: (a) compares collecting data via a ground truth oracle (Local A*) against our “Complete” and “Incomplete” collection methods. (b) plots the “Speed-up” (as measured in nodes expanded) of LoHA* trained on datasets gathered from Local A* calls (true oracle) or our datasets from running a global A* on start-goal problems. (c) We run DE-LoHA* by solving start-goal problems (and collecting data while doing so), and retraining every 5 problems. We see that DE-LoHA* can improve performance from just solving start-goal problems without needing an external oracle after the initial 5 problems.
@@
Closed List as an Obstacle Astute readers may have noticed a possible issue with Theorem 1; although the states are sorted identically, we ignored the effect of the closed list of previously expanded states in the global A*. These states $\in LR(s)$ would not be re-expanded when explored through $s$ in the global search and we can thus view states in the closed list as obstacles. Note this observation has been exploited in previous heuristic search work on grid worlds (Felner, Shperberg, and Buzhish 2021). Interestingly, we found that incorporating the closed list led to no meaningful performance difference so we ignore it.
@@ 4 Experimental Results
Data Efficiency Table 3(a) shows how many nodes expanded are required to collect a single training data point as a function of $K$. We see that running local searches scales poorly and requires 100s of expansions for a single data point for $K>4$. We observe that as $K$ increases, the work for gathering “Complete” via backtracking increases but then saturates as when $K$ is sufficiently high, the global A* mainly adds data points on states on the solution path and few else (as the search rarely fully explores a local minima not on the path). On the flip side, incomplete remains constant regardless of $K$ as all states with any successor expanded becomes a data point (as we have partial data about its $LH(s)$ value), including states that are in local minima.
@@
One concern with using incomplete data by backtracking is that it is noisier data compared to the true residuals collected by local A*. In order to understand the impact of this quality difference, Fig 3(b) plots the performance of LoHA* as a function of the amount of nodes expanded to collect each dataset. The y-axis “Speed-up” is the multiplicative reduction of nodes expanded to find paths across 100 problems between weighted LoHA* and weighted A* with $w=4$ (e.g. 3 represents LoHA* expanding 1/3 the nodes). We see that even though incomplete data (green) contains approximations, we can still effectively use it to learn a heuristic that leads to performance gain in substantially less data. Using a local A* (blue) or only complete data (orange) requires a magnitude more work to get similar performance. We ran an ablation removing downweighing (i.e. not weighting by $\alpha(s,s^{\prime})$) and found that this reduced performance from 3.9x to 2.4x. Note that although LoHA* expands fewer nodes, the model inference introduces a large overhead which results in it taking roughly 2.3 seconds per problem compared to the baseline which takes 0.1 seconds.
@@
Online Performance We evaluate how using our data-efficient framework DE-LoHA* enables learning local heuristics online. We initially start with a small dataset collected by solving 5 start-goal planning problems using global A* and the backtracking logic. This small dataset is used to train an initial local heuristic model. The learnt model is then used with DE-LoHA* to solve another 5 start-goal problems. The data accumulated while solving these problems is accumulated and then used to retrain and improve the model further (the model is retrained after solving every 5 problems). We evaluate online DE-LoHA* run with different $w$ with the corresponding weighted A* baseline. We plot the performance on 50 problems (plotting on just the 5 problems encountered is too noisy). Within 20 encountered start-goal problems (3 iterations of retraining DE-LoHA*), we start to get a non-trivial performance increase. As we continue solving problems and gathering data, improvement increases ($w=16$) or saturates ($w=4,8$). This highlights how DE-LoHA* can perform well from just solving start-goal problems.
@@ 5 Conclusion
Extensions Our data efficient technique can be applied with no or small modifications to other learning methods which require an oracle search method. When learning a cost-to-go heuristic $h(s,s^{\prime})$, existing work just uses states on the solution path with the oracle’s optimal $c(s,s_{goal})$ (Jabbari Arfaee, Zilles, and Holte 2011; Kim and An 2020). We observe how any state $s_{i}$ and an ancestor $s_{i}^{\prime}$ in the oracle’s search tree results in a valid optimal $c(s_{i}^{\prime},s_{i})$ data point that can be used as well. We can thus backtrack from $s_{i}$ to ancestors to collect data. Similarly for learning an expansion delay (Kaur, Chatterjee, and Likhachev 2021), the delays between states $s_{i}^{\prime},s_{i}$ can be used. We hope future work builds on our data efficient framework to decrease computational burden and enable online learning.
Acknowledgements
R.V. would like to thank G.K. for his motivation and advice. This material is partially supported by NSF Grant IIS-2328671.
@@ 6 Procedure Plan($S,E,s_{start},s_{goal},K$)
8       OPEN = $\{s_{start}\}$
9       while OPEN not empty do
10             $s_{min}$ = OPEN.min()
11             if $s_{min}==s_{goal}$ then
@@ 8       OPEN = $\{s_{start}\}$
13
14            Successors = Expand($s_{min}$)
15             $\forall s^{\prime}\in$ Successors, $s^{\prime}$.parent = $s_{min}$triangleright Required anyway to backtrack solution path
16
17             BacktrackDataCollection($s_{min}$)
18             $\forall s^{\prime}\in$ Successors, OPEN.insert($s^{\prime}$)
19
@@ 22 Procedure BacktrackDataCollection($s$)
23       $s_{cur}$ = $s$.parent
24       while $s_{cur}$ != null and $s_{cur}\notin$ CompletedData do
@@
26                   CompletedData[$s_{cur}$] = b ($s_{cur},s$)
27                   Remove $s_{cur}$ from IncompleteData
28
@@ 30                   IncompleteData[$s_{cur}$] = b ($s_{cur},s$)
33
34
35Procedure b($s$, $s^{\prime}$)
36       $c(s,s^{\prime})=s^{\prime}.cost-s.cost$
37       $h_{k}=c(s,s^{\prime})+h_{g}(s^{\prime})-h_{g}(s)$
38       return $h_{k}$
39